# Benchmarking Spiking Neural Network Learning Methods with Varying Locality

### v1

@@ Abstract
Spiking Neural Networks (SNNs), providing more realistic neuronal dynamics, have shown to achieve performance comparable to Artificial Neural Networks (ANNs) in several machine learning tasks. Information is processed as spikes within SNNs in an event-based mechanism that significantly reduces energy consumption. However, training SNNs is challenging due to the non-differentiable nature of the spiking mechanism. Traditional approaches, such as Backpropagation Through Time (BPTT), have shown effectiveness but comes with additional computational and memory costs and are biologically implausible. In contrast, recent works propose alternative learning methods with varying degrees of locality, demonstrating success in classification tasks. In this work, we show that these methods share similarities during the training process, while they present a trade-off between biological plausibility and performance. Further, this research examines the implicitly recurrent nature of SNNs and investigates the influence of addition of explicit recurrence to SNNs. We experimentally prove that the addition of explicit recurrent weights enhances the robustness of SNNs. We also investigate the performance of local learning methods under gradient and non-gradient based adversarial attacks.
@@
Spiking Neural Networks, Local Learning, Training Methods, Centered Kernel Alignment, Fisher Information.I Introduction
@@
Spiking Neural Networks (SNNs) is a type of Artificial Neural Networks (ANNs) that is inspired by biological neurons in the brain [1, 2]. In contrast to traditional ANNs, SNNs use a more biologically realistic model of neuron behavior, where information is processed through the propagation of spikes between neurons in the network [3]. The timing and rate of spikes are influenced by the strength of synaptic connections between neurons, which can be modified through a process called synaptic plasticity [4, 5]. In SNNs, neuronal activities are based on discrete spikes, which occur only when the membrane potential of a neuron exceeds a certain threshold. This event-driven processing operates at significantly lower energy levels than traditional computing architectures [6, 7, 8], as the neuronal and synaptic state updates consume power only when they are actively spiking [9]. Neuromorphic hardware implementations, like Intel’s Loihi [10] and the SpiNNaker project [11], are specifically designed to take advantage of SNNs. The energy efficiency offered by SNNs integrated with neuromorphic hardware provide efficient possibilities for deploying machine learning applications in power-constrained environments, such as mobile devices, wearable technology, and remote sensing systems, among others.
@@
Additionally, since membrane potentials are accumulated over time retaining information from previous time steps, SNNs are inherently recurrent [12]. This characteristic of SNNs allows them to naturally process time-dependent data, rendering them suitable for tasks such as speech recognition or time-series prediction [12, 13, 14, 15, 16]. SNNs with explicit recurrent connections have shown to further augment their performance in speech processing tasks [13].
@@
Although SNNs are demonstrated to provide multiple benefits from algorithms to hardware by leveraging bio-realistic computations, Backpropagation Through Time (BPTT), which is widely used to train SNNs, is not biologically plausible. Firstly, BPTT assumes weight symmetry for the forward and backward passes, which is undesirable in biological neural networks [17]. Secondly, BPTT requires global error propagation, which is inconsistent with the local learning mechanisms observed in biological neural networks [18, 19]. Moreover, the explicit calculation and back-propagation of gradients and error signals in BPTT is not directly observed in the brain [20, 21]. Lastly, BPTT maintains a history of the membrane potential and spike history of each neuron over time. This process can be memory-intensive and computationally expensive [22, 23, 19, 15, 24]. To solve this problem, alternative training methods, such as Spike Timing-Dependent Plasticity (STDP) which is a specific form of Hebbian learning [5, 25] and local learning alternatives to BPTT have been proposed for training SNNs [13, 26, 27, 19, 28]. These methods aim to address the challenges associated with training SNNs while still maintaining biologically plausible learning rules and enabling more efficient and accurate learning in spiking networks.
@@
Based on the the locality of training methods [27], we revisit these alternatives to BPTT. Figure 1 distinguishes SNN training methods based on varying levels of locality. Beginning with BPTT, which is a global learning method, it unrolls neural dynamics over time, and it involves backpropagating errors and learning signals through layers. The connection used for backpropagation is the same weight used for forward propagation. An alternative, named feedback alignment, replaces the symmetric weight connections with random matrices during backward passes. This method removes the need for symmetric weight during learning. E-prop introduces a type of direct feedback alignment that provides higher biological plausibility [15]. In this method, the error calculated at the output layer is propagated directly to each layer without propagating errors layer by layer. To obtain a better performance, e-prop employs symmetric weights in the process of direct propagation. We next consider a more localized method, DECOLLE, in the regime of local error propagation [19]. In this learning rule, errors are propagated at each layer. A cost function is assigned at each layer by a random matrix mapping the output from each layer to the pseudo-targets. In general, all these learning rules provide bio-plausible alternatives to global learning rules.
@@ II-A Performance Evaluation
Recently, the development of novel learning algorithms to train SNNs have gained traction, propelled by SNNs’ unique spiking nature and energy advantage over ANNs. Whether energy efficiency is a compromise of performance remains a question. Ref. [29] explored such a co-design evaluation with respect to ANN performance. Their results indicated that SNNs maintained acceptable accuracy drops on simpler visual tasks like MNIST classification with much better energy efficiency, while in relatively more complicated visual tasks like CIFAR10 classification, the accuracy of SNNs degraded further. Despite this, SNNs evaluated on neuromorphic datasets were illustrated to outperform ANNs. Subsequent progress in SNN training have produced state-of-the-art performance of SNNs in several image recognition benchmarks with recent works indicating highly similar representations learned by residual SNNs and ANNs using representational similarity measurements [30].
@@
Moreover, the inherently recurrent nature of SNNs resembles the operation of recurrent neural networks (RNNs) in terms of temporal and spatial neural dynamics. Inspired by such characteristics, prior work has compared SNNs and RNNs, highlighting the superior performance of SNNs over RNNs on neuromorphic datasets [31]. Additionally, the work investigated the influence of temporal resolution, leakage and reset mechanisms on the performance of SNNs, illustrating that the exclusion of leakage and reset mechanisms negatively impacted the performance, and the increment of temporal resolution was not necessarily useful. In these benchmarking works, SNNs were trained with the BPTT training method, where surrogate functions were used to overcome the non-differentiability of spiking neurons. To quantitatively analyze the impact of surrogate functions on the performance of SNNs, experimental results from previous work have demonstrated that surrogate functions help recurrent SNNs outperform conventional RNNs [14]. Additionally, this work examined the advantage of adaptive neurons [13], where thresholds are updated based on firing rates, and illustrated that recurrent SNNs with adaptive neurons gain performance advantages over SNNs with fixed thresholds. Mixed usage of both types of neurons was shown to improve the performance of SNNs over LSTM architectures [13].
@@
In conclusion, prior works have mainly targeted benchmarking comparable performance of SNNs and ANNs. Some works have investigated the influence of various SNN-specific control knobs like leakage, temporal dynamics, among others on SNN performance. However, the majority of benchmarking efforts optimize SNNs via the BPTT algorithm, thereby not providing insights related to emerging SNN learning schemes offering local learning possibilities. In contrast to prior works, we perform an extensive benchmarking analysis with a complementary perspective where we investigate the performance of SNNs with regards to learning methods in terms of accuracy and robustness with both implicit and explicit recurrent architectures.
@@ II-B Adversarial Attacks
The adversarial robustness of SNNs has been proven by various researchers. Sharmin et al. [32] explored gradient-based attacks and its variants on both the ANN-converted SNN and SNN trained from scratch and found SNN’s robustness against adversarial attacks. Similarly, Liang et al. [33] conducted untargeted gradient attacks, and found that SNNs require more perturbations for successful attacks. Marchisio et al. [34] considered a deep Spiking Deep Belief Network (SDBN) using noise attacks with the same motive of backdoor trigger attacks [35] and reached similar conclusions. However, their studies were either focused on a single training method or relied entirely on the gradient. However, while at its exploration stage, many recent SNN training methods look for alternatives to using gradient information due to high computational costs [24, 36]. In this work, we aim to bridge this gap by revisiting adversarial attacks on SNNs and provide a benchmarking analysis for non-gradient poisoning attacks as well. The benchmarking effort is evaluated corresponding to various SNN training paradigms for future researchers to make more informed design choices.
@@ III Methods
Sections A-B discusses SNN computational models and learning algorithms considered in this work while sections C-D introduces empirical tools used to understand the learning dynamics of various SNN topologies and learning methods.
@@ III-A Spiking Neurons
In this study, we use SNNs with Leaky-Integrate-and-Fire (LIF) neurons [14, 27]. The membrane potential and synaptic currents for a layer at time step in a discrete time setting is expressed as:
Here, represents the weight connection between the current layer and the previous layer , and represents the recurrent weight connection of current layer. denotes the output of the previous layer at time step . The parameters and correspond to the synaptic and membrane decay rate respectively. Each neuron maintains a membrane potential , and if it surpasses the threshold of the spiking neuron, a spike will be fired. The membrane potential is subtracted by the threshold after each firing event, which is captured by the refractory period .
@@ III-B1 Backpropagation Through Time (BPTT)
where, is the total number of time steps, is the learning rate, and the gradient of the loss function with respect to the weight in layer at time step can be expressed as:
where, represents the error term for neuron in layer at time step , and denotes the activation of neuron in layer at time step . The error term can be computed recursively as:
where, is the number of neurons in layer , is the weight connecting neuron in layer to neuron in layer at time step . represents the derivative of the surrogate function with respect to the pre-activation value for neuron in layer at time step .
@@ III-B2 E-prop
Recent studies of error propagation methods in SNNs propose a new method called e-prop [15]. Compared with BPTT, e-prop offers higher biological plausibility, as it is inspired by the concept of synaptic eligibility traces observed in neuroscience [38], which weighs the temporal differences of presynaptic and postsynaptic neuronal activations. While BPTT relies on a global learning rule that requires error information to be propagated backward through the network between layers, e-prop aims to derive a local learning rule, in which the error signals are propagated directly from the output layer to hidden units. The local learning rule in e-prop makes it more suitable for neuromorphic hardware implementations, where global communication can be a significant challenge.
In the e-prop method, the weight updates are based on the product of eligibility traces and learning signals. The eligibility trace for a synapse connecting neuron to neuron at time step can be computed by:
where, represents the pre-activation of neuron at time step , and is the weight of the synapse connecting neuron to neuron . The eligibility trace captures the recent history of a synapse’s contribution to the post-synaptic neuron’s activation. The learning signal for neuron at time step is calculated as:
where, is the loss function and is the activation of neuron at time step . The learning signal quantifies how the activation of neuron at time step contributes to the overall loss. The weight update for a synapse connecting neuron to neuron can be expressed as:
where, is the learning rate and is the total number of time steps over which the network is operated.
@@ III-B3 DECOLLE
Deep Continuous Local Learning (DECOLLE) is a biologically plausible online learning algorithm for training SNNs [19]. DECOLLE combines the principles of local learning with deep architectures, enabling the training of SNNs in a layer-by-layer manner without the need for backpropagation or weight transport. In DECOLLE, each layer in the network learns features using local information, which allows the network to adapt online without relying on global error signals. This is achieved by attaching an auxiliary cost function to the random readouts at each layer. The random readouts at layer is calculated as:
where, is a fixed random matrix for layer and is the activation from neuron that connects to neuron . Then, the loss function is calculated depending on the differences between readouts and pseudo-target . Based on this, weight updates at each layer are determined using a local learning rule:
where, represents the learning rate, denotes the surrogate gradient of the non-differentiable step function, refers to the traces of the membrane driven solely by incoming spikes, with a constant factor captures the refractory dynamics depending on the spiking history of the neuron. DECOLLE ignores the spiking history dependencies, considering to have a negligible impact on the membrane potential. Consequently, the synaptic weight updates become:
Note that the locality of this algorithm is maintained by setting all non-local gradients to zero, as expressed by the equation:
This constraint ensures that the learning process relies exclusively on local information, adhering to the principles of biologically plausible learning rules.
III-C Fisher Information
@@
As SNNs rely on accumulating input data over multiple time steps to predict class probabilities, the Fisher Information Matrix (FIM) is accumulated over the time dimension to analyze the information dynamics across each time step. The amount of Fisher Information present in SNNs at a given time step is the sum of information across all previous time steps, from 1 to .
Here, represents the index of the time step. Direct calculation on the Fisher Information Matrix (FIM) is computationally intensive, due to the large number of parameters in an SNN. Consequently, computing the trace of the FIM is a more efficient alternative [40]. Given a dataset of training samples, we can compute the Fisher Information for each layer at time as follows:
@@
III-D Centered Kernel Alignment
Centered Kernel Alignment (CKA) is used to quantify the similarity between representations in two arbitrary layers through normalized Hilbert-Schmidt Independence Criterion (HSIC) [41, 30]. In this work, the CKA calculation is adopted from Li et al. [30].

### v2

@@ Abstract
Spiking Neural Networks (SNNs), providing more realistic neuronal dynamics, have been shown to achieve performance comparable to Artificial Neural Networks (ANNs) in several machine learning tasks. Information is processed as spikes within SNNs in an event-based mechanism that significantly reduces energy consumption. However, training SNNs is challenging due to the non-differentiable nature of the spiking mechanism. Traditional approaches, such as Backpropagation Through Time (BPTT), have shown effectiveness but come with additional computational and memory costs and are biologically implausible. In contrast, recent works propose alternative learning methods with varying degrees of locality, demonstrating success in classification tasks. In this work, we show that these methods share similarities during the training process, while they present a trade-off between biological plausibility and performance. Further, given the implicitly recurrent nature of SNNs, this research investigates the influence of the addition of explicit recurrence to SNNs. We experimentally prove that the addition of explicit recurrent weights enhances the robustness of SNNs. We also investigate the performance of local learning methods under gradient and non-gradient-based adversarial attacks.
@@
Spiking Neural Networks, Local Learning, Training Methods, Feedback Alignment, Direct Feedback Alignment, Adversarial Attack, Backdoor Attack.I Introduction
@@
Spiking Neural Networks (SNNs) are a type of Artificial Neural Networks (ANNs) that are inspired by biological neurons in the brain [1, 2]. In contrast to traditional ANNs, SNNs use a more biologically realistic model of neuron behavior, where information is processed through the propagation of spikes between neurons in the network [3]. The timing and rate of spikes are influenced by the strength of synaptic connections between neurons, which can be modified through a process called synaptic plasticity [4, 5]. In SNNs, neuronal activities are based on discrete spikes, which occur only when the membrane potential of a neuron exceeds a certain threshold. This event-driven processing operates at significantly lower energy levels than traditional computing architectures [6, 7, 8], as the neuronal and synaptic state updates consume power only when they are actively spiking [9]. Neuromorphic hardware implementations, like Intel’s Loihi [10] and the SpiNNaker project [11], are specifically designed to take advantage of SNNs. The energy efficiency offered by SNNs integrated with neuromorphic hardware provides efficient possibilities for deploying machine learning applications in power-constrained environments, such as mobile devices, wearable technology, and remote sensing systems, among others.
@@
Additionally, since membrane potentials are accumulated over time retaining information from previous time steps, SNNs are inherently recurrent [12]. This characteristic of SNNs allows them to naturally process time-dependent data, rendering them suitable for tasks such as speech recognition or time-series prediction [12, 13, 14, 15, 16]. SNNs with explicit recurrent connections have been shown to further augment their performance in speech processing tasks [13].
@@
Although SNNs are demonstrated to provide multiple benefits from algorithms to hardware by leveraging bio-realistic computations, Backpropagation Through Time (BPTT), which is widely used to train SNNs, is not biologically plausible. Firstly, BPTT assumes weight symmetry for the forward and backward passes, which is undesirable in biological neural networks [17]. Secondly, BPTT requires global error propagation, which is inconsistent with the local learning mechanisms observed in biological neural networks [18, 19]. Moreover, the explicit calculation and back-propagation of gradients and error signals in BPTT are not directly observed in the brain [20, 21]. Lastly, BPTT maintains a history of the membrane potential and spike history of each neuron over time. This process can be memory-intensive and computationally expensive [22, 23, 19, 15, 24]. To solve this problem, local learning alternatives to BPTT have been proposed for training SNNs [13, 25, 26, 19, 27]. These methods aim to address the challenges associated with training SNNs while still maintaining biologically plausible learning rules and enabling more efficient and accurate learning in spiking networks.
@@
Based on the locality of training methods [26], we revisit these alternatives to BPTT. Figure 1 distinguishes SNN training methods based on varying levels of locality. Beginning with BPTT, which is a global learning method, it unrolls neural dynamics over time, and it involves backpropagating errors and learning signals through layers. The connection used for backpropagation is the same weight used for forward propagation. An alternative, named feedback alignment, replaces the symmetric weight connections with random matrices during backward passes. This method removes the need for symmetric weights during learning. E-prop introduces a type of direct feedback alignment that provides higher biological plausibility [15]. In this method, the error calculated at the output layer is propagated directly to each layer without propagating errors layer by layer. To obtain better performance, e-prop employs symmetric weights in the process of direct propagation. We next consider a more localized method, DECOLLE, in the regime of local error propagation [19]. In this learning rule, errors are propagated at each layer. A cost function is assigned at each layer by a random matrix mapping the output from each layer to the pseudo-targets. In general, all these learning rules provide bio-plausible alternatives to global learning rules.
@@ II-A Performance Evaluation
Recently, the development of novel learning algorithms to train SNNs has gained traction, propelled by SNNs’ unique spiking nature and energy advantage over ANNs. Whether energy efficiency is a compromise of performance remains a question. Ref. [28] explored such a co-design evaluation with respect to ANN performance. Their results indicated that SNNs maintained acceptable accuracy drops on simpler visual tasks like MNIST classification with much better energy efficiency, while in relatively more complicated visual tasks like CIFAR10 classification, the accuracy of SNNs degraded further. Despite this, SNNs evaluated on neuromorphic datasets were illustrated to outperform ANNs. Subsequent progress in SNN training has produced state-of-the-art performance of SNNs in several image recognition benchmarks, with recent works indicating highly similar representations learned by residual SNNs and ANNs using representational similarity measurements [29].
@@
Moreover, the inherently recurrent nature of SNNs resembles the operation of recurrent neural networks (RNNs) in terms of temporal and spatial neural dynamics. Inspired by such characteristics, prior work has compared SNNs and RNNs, highlighting the superior performance of SNNs over RNNs on neuromorphic datasets [30]. Additionally, the work investigated the influence of temporal resolution, leakage, and reset mechanisms on the performance of SNNs, illustrating that the exclusion of leakage and reset mechanisms negatively impacted the performance, and the increment of temporal resolution was not necessarily useful. In these benchmarking works, SNNs were trained with the BPTT training method, where surrogate functions were used to overcome the non-differentiability of spiking neurons. To quantitatively analyze the impact of surrogate functions on the performance of SNNs, experimental results from previous work have demonstrated that surrogate functions help recurrent SNNs outperform conventional RNNs, and the choice of surrogate functions exhibits negligible impact on performance [14]. Additionally, this work examined the advantage of adaptive neurons [13], where thresholds are updated based on firing rates, and illustrated that recurrent SNNs with adaptive neurons gain performance advantages over SNNs with fixed thresholds. Mixed usage of both types of neurons was shown to improve the performance of SNNs over LSTM architectures [13].
@@
In conclusion, prior works have mainly targeted benchmarking comparable performance of SNNs and ANNs. Some works have investigated the influence of various SNN-specific control knobs like leakage, temporal dynamics, among others, on SNN performance. However, the majority of benchmarking efforts optimize SNNs via the BPTT algorithm, thereby not providing insights related to local learning methods. In contrast to prior works, we perform an extensive benchmarking analysis with a complementary perspective where we investigate the performance of SNNs with regard to local learning methods in terms of accuracy and robustness with both implicit and explicit recurrent architectures.
@@ II-B Adversarial Attacks
The adversarial robustness of SNNs has been proven by various researchers. Sharmin et al. [31] explored gradient-based attacks and their variants on both the ANN-converted SNN and SNN trained from scratch and found SNN’s robustness against adversarial attacks. Similarly, Liang et al. [32] conducted untargeted gradient attacks and found that SNNs require more perturbations for successful attacks. Marchisio et al. [33] considered a deep Spiking Deep Belief Network (SDBN) using noise attacks with the same motive of backdoor trigger attacks [34] and reached similar conclusions. However, their studies were either focused on a single training method or relied entirely on the gradient. While at its exploration stage, many recent SNN training methods look for alternatives to using local training methods [24, 35]. In this work, we aim to bridge this gap by revisiting adversarial attacks on SNNs trained by local learning methods and provide a benchmarking analysis for non-gradient poisoning attacks as well. Although adversarial training strategies strengthen machine learning models against adversarial inputs in the domain of SNNs [36, 37, 38], these mechanisms are out of the scope of this study.
The primary contributions of this work are the following:
We extensively benchmark SNNs’ performance, concentrating on learning methods such as BPTT, e-prop, and DECOLLE applied to both implicit and explicit recurrent architectures. We highlight the trade-off between biological plausibility and performance, as well as the enhancement of accuracy achieved by incorporating additional weights in explicit recurrent SNNs.
We perform an in-depth analysis to reveal that training methods with similar locality share representational similarities. Additionally, we find recurrent and linear weights have comparable importance, with each linear layer’s significance growing from input to output.
This study offers insights into the robustness of local learning methods by revisiting adversarial attacks for both gradient-based and non-gradient poisoning attacks. SNNs trained with local learning methods are more robust than those using global learning methods for gradient-based attacks. Analysis shows that explicit recurrent SNNs have greater robustness, supported by the Centered Kernel Alignment (CKA) metric.
@@ III Methods
Sections A-B discuss SNN computational models and learning algorithms considered in this work. Section C analyzes the computational complexity of learning methods. Sections D-E introduce empirical tools used to understand the learning dynamics of various SNN topologies and learning methods.
@@ III-A Spiking Neurons
In this study, we use SNNs with Leaky-Integrate-and-Fire (LIF) neurons [14, 26]. The membrane potential and synaptic currents for a layer at time step in a discrete time setting are expressed as:
Here, represents the weight connection between the current layer and the previous layer , and represents the recurrent weight connection of the current layer. denotes the activation of the previous layer at time step . The parameters and correspond to the synaptic and membrane decay rates respectively. Each neuron maintains a membrane potential , and if it surpasses the threshold of the spiking neuron, a spike will be fired. The membrane potential is subtracted by the threshold after each firing event, which is captured by the refractory period .
@@ III-B1 Backpropagation Through Time (BPTT)
where, is the total number of time steps, is the learning rate, and is the loss function. At time step , represents the weight connections between neurons and , is the activation of neuron , and approximates the derivative of the activation function with respect to the pre-activation value for neuron with a surrogate function.
@@ III-B2 E-prop
Recent studies of error propagation methods in SNNs propose a new method called e-prop [15]. Compared with BPTT, e-prop offers higher biological plausibility, as it is inspired by the concept of synaptic eligibility traces observed in neuroscience [40], which weigh the temporal differences of presynaptic and postsynaptic neuronal activations. While BPTT relies on a global learning rule that requires error information to be propagated backward through the network between layers, e-prop aims to derive a local learning rule, in which the error signals are propagated directly from the output layer to hidden units. The local learning rule in e-prop makes it more suitable for neuromorphic hardware implementations.
In the e-prop method, the weight updates are based on the product of eligibility traces and learning signals. The eligibility trace for a synapse connecting neuron to neuron at time step can be computed recursively by eligibility vector :
The eligibility trace captures the recent history of a synapse’s contribution to the post-synaptic neuron’s activation. The learning signal for neuron at time step , following direct feedback alignment, routes the error calculated at the output layer to the current neuron:
where is a fixed random matrix and captures the loss calculated at the output layer. The weight update can be expressed as:
@@ III-B3 DECOLLE
Deep Continuous Local Learning (DECOLLE) is a biologically plausible online learning algorithm for training SNNs [19]. DECOLLE combines the principles of local learning with deep architectures, enabling the training of SNNs in a layer-by-layer manner without the need for backpropagation or weight transport. In DECOLLE, each layer in the network learns features using local information, which allows the network to adapt online without relying on global error signals. This is achieved by attaching an auxiliary cost function to the random readouts at each layer. The layer-specific random readouts at time are calculated as:
where is a fixed random matrix for layer . Then, the loss is calculated as the sum of layerwise differences between readouts and pseudo-target . Based on this, weight updates at each layer for a specific time step are determined using a local learning rule:
where refers to the traces of the membrane potential of neuron driven solely by incoming spikes, with a constant factor that captures the refractory dynamics depending on the spiking history of the neuron . DECOLLE ignores the spiking history dependencies, considering to have a negligible impact on the membrane potential. Consequently, the synaptic weight updates become:
Ignoring refractory dynamics ensures that the learning process relies exclusively on local information, adhering to the principles of biologically plausible learning rules.
III-C Computational Advantages of Local Learning Methods
Method | Space | Time |
BPTT | ||
e-prop | ||
DECOLLE |
Table I provides the computational complexity for the aforementioned training methods. BPTT performs backpropagation on unrolled neuron states across time , resulting in a space complexity of and a time complexity of , as it multiplies error signals with gradients computed in the current layer with neurons over . The e-prop learning rule maintains number of eligibility traces (Equation 3) in forward manner for each synaptic connection, leading to a space complexity of [15]. During the gradient computation, the error signals of size are multiplied by the eligibility traces (Equation 5) with a time complexity of . DECOLLE performs online weight updates at each time step [19], using available neuron states without storing, resulting in a space complexity of . During the weight gradient computation, local errors of size are computed (Equation 6) using multiplication operations for neurons. Then these local errors are multiplied by the number of inputs , yielding a total time complexity of .
III-D Fisher information
@@
As SNNs rely on accumulating input data over multiple time steps to predict class probabilities, the amount of Fisher information Matrix (FIM) accumulated in SNNs at a given time step is the sum of information across all previous time steps, from 1 to .
Here, is input image sampled from data distribution , is the output variable, and represents the index of the time step. Direct calculation on the Fisher information Matrix (FIM) is computationally expensive, due to the large number of parameters in an SNN. Consequently, computing the trace of the FIM is a more efficient alternative [42]. Given a dataset of training samples, we can compute the Fisher information at time as follows:
@@
III-E Centered Kernel Alignment (CKA)
CKA is used to quantify the similarity between representations in two arbitrary layers through the normalized Hilbert-Schmidt Independence Criterion (HSIC) [43, 29]. In this work, the CKA calculation is adopted from Li et al. [29].